A certain college student and his roommate just had to listen to Pink Floyd’s The Wall
from start to finish. This sort of endeavor entails frequent trips to the stereo to turn
up the volume, particularly when it’s fun to watch the speakers shake during great
songs with awesome guitar solos.2
If the speakers could handle 50 W of power and had
a resistance of 4.0 Ω, what amperage fuse should have been placed in the circuit to
prevent blowing out the speakers?
1
Expert's answer
2017-02-28T12:12:05-0500
I = sqrt(P/R) = sqrt(50/4) = 3.5 A